# Visual Hand Gesture Recognition with Deep Learning: A Comprehensive Review of Methods, Datasets, Challenges and Future Research Directions

### v2

@@ 3.5 Camera Point of View
Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer (EgoGesture_DATASET). The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design. Three primary viewpoints are considered (see Fig. 5):
@@
(a) First-person view (egocentric) (EgoGesture_DATASET)
(b) Second-person view (Lsa64argentiniansignlanguage)
(c) Third-person view (Auslan-Daily)
@@ 3.7 Learning Strategy
    Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester (Jester-Dataset)) and subsequently fine-tuned for specific downstream VHGR tasks (pretrained_on_jester_3dDenseNet_LSTM).
@@
    Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries (Fangyun_Wei_ICCV).
@@
    Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency (VAC-CSLR; C2ST; NEURIPS2022_6cd3ac24_HEATMAPS).
@@
    Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374), or to improve the discriminative capacity of learned representations (MASA; Weichao-Zhao).
@@
    Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization (UNISIGNli2025).
@@
    Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations (KEYFRAMES_DRL_SKELETON).
@@
    Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance (ZERO-SHOT-SLR; ZS-GR; FEW-SHOT-BILGE2024110374).
@@
    Adversarial Learning, used as a data augmentation strategy (alimisis2025advances) to improve robustness, as demonstrated in adversarial video synthesis networks (ADVERSARIAL_DATA_AUGM).
@@
    Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) (BEST; SignBERT+; SignBERT_ICCV).
@@ 3.8 Processing
Existing methods for isolated and continuous dynamic HGR can be further categorized based on whether the models have full access to the entire gesture(s) sequence or not. According to (NVIDIA_GESTURE_R3DCNN_Molchanov), two types of processing strategies apply:
@@
    Online (real-time) processing: Online HGR models perform gesture recognition in a stream of input frames. In other words, these models take as input the gesture sequence recorded until a specific time. Thus, online gesture recognition is suitable for real-time applications, as the gesture is recognized before the retraction part, reducing the latency. Sliding windows and gesture detector/classifier pipelines are often employed (IPN-HAND; survey-emporio2025continuous).
@@
    Offline: Offline processing means that the model treats the entire gesture sequence. This approach is limited to applications that do not require real-time recognition. Most of the existing methods for VHGR fall into this category, such as those using BiLSTM for temporal modeling (VAC-CSLR).
@@ 4.1 Sign Language Understanding (SLU)
For SLU, the authors in (G-CNN-SHARMA2021115657) developed Gesture-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset. The authors in (STN-ND-Net) proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2DCNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images. The success of vision transformer (ViT) motivated authors of SIGNFORMER (SIGNFORMER) to adopt it for ISL recognition, while authors in (Jungpil-Shin) leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion. A YOLO variant, namely YOLOv4-CSP(YOLOv4-CSP) has also been developed, to simultaneous detect and identify Turkish Sign Language gestures.
@@ 4.2 Human-Computer Interaction (HCI)
For gesture-based HCI the authors in (LHGR-Net) proposed a lightweight 2DCNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM. The authors in (2DPSTPP) proposed 2DPSTPP-Net, a 2DCNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios. The authors in (WEI_HAN_YOLOV5_COPULA) modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes. In (Mohammad_Mahmudul_static_hgr), a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification. In (tuan_linh_pham_dang2023lightweight), a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2DCNN like MobileNet or EfficientNet. (FGDSNet) introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration. The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
@@ 4.3 Human-Robot Interaction (HRI)
HRI, particularly gesture-based mobile robot navigation, has attracted research interest. URGR (ULTRA-RANGE-UGV-Bamani) is a three-stage approach using YOLOv3 for human cropping, High-Quality Network (HQ-Net) for super-resolution to address long-range scenarios, and a classifier, Graph-Vision Transformer (GViT), for gesture recognition. The HQ-Net comprises convolutional and deconvolutional pathways. Features extracted by Canny Edge algorithm and a convolutional module followed by self-attention are fused with the latent representation. GViT utilizes two GCN layers that capture fine-grained details followed by a modified ViT. This setting enables effective recognition of gestures in a distance of up to 25m for UGV guidance. Similarly, (DRCAM_TIM) introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections between the cascaded blocks. Because of the attention mechanisms, the model focuses effectively on the most informative regions of the image.
@@ 4.4 General Purpose
General-purpose models have also been developed. EDenseNet (EDenseNet) extends DenseNet by adding convolutional layers between dense blocks. SpAtNet (SpAtNet) is a lightweight 2DCNN with three streams: two Multi-scale Attentive Feature Fusion (MAFF) modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification. This approach, employing multiple kernel sizes, increases accuracy by capturing various levels of granularity. MMFANet (MMFANet) performs gesture detection and recognition using a 2DCNN backbone followed by the proposed feature aggregation pyramid network (FAPN) and task-specific regressor and classifier branches. The FAPN module aims to extract multiscale features through upsampling/downsampling operations and fusion of adjacent feature maps, while reducing the computational complexity by utilizing novel expanded convolutional blocks with different dilation rates combined with attention mechanisms that eliminate redundant information.
@@ 4.5 Performance of SHGR Methods
Most existing methods for Static Hand Gesture Recognition (SHGR) are evaluated on different or custom datasets, which prevents a direct comparison of their performance. Table 2 presents the accuracy of current state-of-the-art SHGR approaches. Despite their high accuracy and low parameter count—suggesting that SHGR might be a largely solved problem—the datasets used (ASL-FS (ASL-FS-2011-CVPRW), HGR1 (HGR-1-2024-PATTER-RECOGNITION), NUS-I (NUS-I-2010), NUS-II (NUS-II-2013), and OUHANDS (ouhands)) remain relatively small and lack diversity. Consequently, model performance would likely degrade in real-world, uncontrolled environments. Among the existing methods, LHGR-Net (LHGR-Net) and FGDSNet (FGDSNet) demonstrate strong results on the OUHANDS dataset with lightweight architectures, making them suitable for gesture-based human–computer interaction. On the other hand, EDenseNet (EDenseNet) achieves high accuracy on ASL-FS but at a higher computational cost, making it better suited for static sign language recognition. Finally, applying background removal through hand segmentation does not appear to significantly improve accuracy compared to end-to-end learning approaches.
@@
FGDSNet (FGDSNet) * 2024 RGB 1.20M - 94.80% - 99.80% 97.80%
SpAtNet (SpAtNet) 2024 RGB 3.8M - 93.33% - - -
(tuan_linh_pham_dang2023lightweight) * 2023 RGB 5.34M - 98.15% - - 89.20%
LHGR-Net (LHGR-Net) 2023 RGB 1.81M - 93.15% - - 98.57%
2DPSTPP-Net (2DPSTPP) 2023 RGB 19.01M - - 100% 99.61% -
DRCAM (DRCAM_TIM) * 2023 Depth - 93.44% - - - -
EDenseNet (EDenseNet) 2021 RGB 7.77M 99.99% - - 99.30% -
@@ 5.1.1 Spatiotemporal
In (2plus1-D-SLR_and_NCSL), the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution). This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment. To enhance temporal modeling, authors in (Snehal_Bharti_et_al) proposed a trainable key frame selection method based on error correction. VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3DCNN for classification.
In (ShiqiWang_et_al), a dual-stage multi-stream pipeline for CSL recognition was developed. An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands. The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP. Authors in (TS3C-Net) introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation. Finally, (SeST) addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module. The fused features are passed through fully connected and softmax layers for classification.
@@ 5.1.2 Spatial + Temporal
In (DAVID_RICHARD_TOM_HAX), a method tailored for HCI tasks using only depth data was proposed. The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling. For ISLR, authors in (GoogleNet-BiLSTM) introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers. They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in (TIM-SLR). The model uses a 2DCNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation. Finally, the Swin Transformer presented in (Full_Transformer_Network) is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
@@ 5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
EgoFormer (EgoFormer_METHOD_EgoDriving_DATASET) is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving. EgoFormer combines a custom 3DCNN with a transformer module and an MLP for final classification. For long-range gesture-based UGV control, authors in (SlowFast-Transformer) proposed a pipeline where a lightweight 2DCNN extracts per-frame features, which are used to select key frames. After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
To address data scarcity in sign language understanding, (ZERO-SHOT-SLR) proposed a zero-shot learning framework. Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes. This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
@@ 5.1.4 Other
Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR). For example, the method proposed in (HAMZAH_LUQMAN) for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet. A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM. The outputs of both streams are concatenated and refined using a custom shallow 2DCNN. A different approach was presented in (Ensemble-InceptionV4), where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
In the context of pose-based ISLR, (Tianyu_Liu) proposed a dual-stream model with keyframe selection. A 2DCNN and attention mechanism filter out less informative frames. The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame. The fused features are used for final classification. To improve signer-independence and reduce overfitting, the SwC GR-MMixer (SwC-GR-MMixer), another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.

### v3

@@ 3.5 Camera Point of View
Complementing the previous criteria on gesture content and input modalities, this dimension focuses on the spatial relationship between the camera and the performer (220). The camera’s perspective influences both the visual characteristics of the gesture and the recognition model design. Three primary viewpoints are considered (see Fig. 5):
@@
(a) First-person view (egocentric) (220)
(b) Second-person view (167)
(c) Third-person view (177)
@@ 3.7 Learning Strategy
    Transfer Learning, where models are pretrained on large-scale datasets (e.g., Jester (131)) and subsequently fine-tuned for specific downstream VHGR tasks (125).
@@
    Weakly Supervised Learning, often used in Continuous Hand Gesture Recognition (CHGR), where temporal annotations are coarse or unaligned, lacking exact gesture boundaries (207).
@@
    Knowledge Distillation, in which one part of the network (often a teacher model) transfers knowledge to another (student model) to enhance feature extraction and efficiency (136; 217; 29).
@@
    Contrastive Learning, used either in a CLIP-based framework to support zero-shot/few-shot recognition by leveraging textual descriptions (23; 164; 24), or to improve the discriminative capacity of learned representations (223; 225).
@@
    Multi-task Learning, where a single model is optimized to perform multiple related tasks simultaneously, encouraging knowledge sharing and improved generalization (113).
@@
    Reinforcement Learning, applied to optimize decision-making processes, such as keyframe selection, by rewarding effective representations (51).
@@
    Few/Zero-shot Learning, which enable the model to generalize to new gestures or languages with limited or no annotated examples, often leveraging semantic or textual guidance (23; 164; 24).
@@
    Adversarial Learning, used as a data augmentation strategy (13) to improve robustness, as demonstrated in adversarial video synthesis networks (142).
@@
    Masked Reconstruction Pretraining, inspired by BERT-style architectures, in which models are pretrained to reconstruct masked skeleton sequences, injecting structural prior knowledge (e.g., hand topology) (224; 77; 78).
@@ 3.8 Processing
Existing methods for isolated and continuous dynamic HGR can be further categorized based on whether the models have full access to the entire gesture(s) sequence or not. According to (138), two types of processing strategies apply:
@@
    Online (real-time) processing: Online HGR models perform gesture recognition in a stream of input frames. In other words, these models take as input the gesture sequence recorded until a specific time. Thus, online gesture recognition is suitable for real-time applications, as the gesture is recognized before the retraction part, reducing the latency. Sliding windows and gesture detector/classifier pipelines are often employed (20; 47).
@@
    Offline: Offline processing means that the model treats the entire gesture sequence. This approach is limited to applications that do not require real-time recognition. Most of the existing methods for VHGR fall into this category, such as those using BiLSTM for temporal modeling (136).
@@ 4.1 Sign Language Understanding (SLU)
For SLU, the authors in (176) developed Gesture-CNN, a shallow 2DCNN to recognize Indian Sign Language (ISL) signs and slightly outperformed various versions of VGG on a self-collected dataset. The authors in (60) proposed the STN-ND-Net, a combination of a Spatial Transformer Network (STN) that improves robustness to spatial transformations, with a 2DCNN backbone that uses deconvolution operations to reduce the pixel-wise and channel-wise correlations and eliminates the information redundancy of colored images. The success of vision transformer (ViT) motivated authors of SIGNFORMER (105) to adopt it for ISL recognition, while authors in (180) leveraged both skeleton and RGB data in their dual stream feature extractor and an MLP to perform intermediate fusion. A YOLO variant, namely YOLOv4-CSP(8) has also been developed, to simultaneous detect and identify Turkish Sign Language gestures.
@@ 4.2 Human-Computer Interaction (HCI)
For gesture-based HCI the authors in (218) proposed a lightweight 2DCNN, named LHGR-Net, which combines a Multiscale Structure (MSS) that aims to make the model scale-invariant via atrous convolution with a Lightweight Attention Structure (LAS) based on CBAM. The authors in (89) proposed 2DPSTPP-Net, a 2DCNN with four stacked convolutional blocks whose features are pooled using max pooling and a novel method, then concatenated to form the final representation, tailored for HCI scenarios. The authors in (67) modified the YOLOv5 to make it more lightweight and developed a model that can be used for gesture-based user interfaces in smart sports stadiums, improving the experience of spectators and athletes. In (9), a pipeline for egocentric HGR and fingertip localization uses YOLOv2 for hand detection, followed by VGG16 for feature extraction and gesture classification. In (36), a lightweight HCI pipeline segments the hand using DeepLabV3+ or Unet, then classifies the cropped image with a 2DCNN like MobileNet or EfficientNet. (229) introduced FGDSNet, a two-stage approach with dual-stream segmentation (local detail and semantic branches) and multi-level fusion for feature integration. The segmentation mask and original image are fed into a dual-stream classification module, where mid-level fusion combines features for final classification via an MLP.
@@ 4.3 Human-Robot Interaction (HRI)
HRI, particularly gesture-based mobile robot navigation, has attracted research interest. URGR (18) is a three-stage approach using YOLOv3 for human cropping, High-Quality Network (HQ-Net) for super-resolution to address long-range scenarios, and a classifier, Graph-Vision Transformer (GViT), for gesture recognition. The HQ-Net comprises convolutional and deconvolutional pathways. Features extracted by Canny Edge algorithm and a convolutional module followed by self-attention are fused with the latent representation. GViT utilizes two GCN layers that capture fine-grained details followed by a modified ViT. This setting enables effective recognition of gestures in a distance of up to 25m for UGV guidance. Similarly, (171) introduced a two-stage method based on depth data: hand segmentation followed by gesture classification using DRCAM, a 2DCNN with stacked Residual Channel Attention Modules (RCAMs) and dense interconnections between the cascaded blocks. Because of the attention mechanisms, the model focuses effectively on the most informative regions of the image.
@@ 4.4 General Purpose
General-purpose models have also been developed. EDenseNet (189) extends DenseNet by adding convolutional layers between dense blocks. SpAtNet (22) is a lightweight 2DCNN with three streams: two Multi-scale Attentive Feature Fusion (MAFF) modules for coarse-grained features and one Interleaved Module for high-level representations; their fused features are fed to an MLP for classification. This approach, employing multiple kernel sizes, increases accuracy by capturing various levels of granularity. MMFANet (114) performs gesture detection and recognition using a 2DCNN backbone followed by the proposed feature aggregation pyramid network (FAPN) and task-specific regressor and classifier branches. The FAPN module aims to extract multiscale features through upsampling/downsampling operations and fusion of adjacent feature maps, while reducing the computational complexity by utilizing novel expanded convolutional blocks with different dilation rates combined with attention mechanisms that eliminate redundant information.
@@ 4.5 Performance of SHGR Methods
Most existing methods for Static Hand Gesture Recognition (SHGR) are evaluated on different or custom datasets, which prevents a direct comparison of their performance. Table 2 presents the accuracy of current state-of-the-art SHGR approaches. Despite their high accuracy and low parameter count—suggesting that SHGR might be a largely solved problem—the datasets used (ASL-FS (154), HGR1 (98), NUS-I (106), NUS-II (153), and OUHANDS (132)) remain relatively small and lack diversity. Consequently, model performance would likely degrade in real-world, uncontrolled environments. Among the existing methods, LHGR-Net (218) and FGDSNet (229) demonstrate strong results on the OUHANDS dataset with lightweight architectures, making them suitable for gesture-based human–computer interaction. On the other hand, EDenseNet (189) achieves high accuracy on ASL-FS but at a higher computational cost, making it better suited for static sign language recognition. Finally, applying background removal through hand segmentation does not appear to significantly improve accuracy compared to end-to-end learning approaches.
@@
FGDSNet (229) * 2024 RGB 1.20M - 94.80% - 99.80% 97.80%
SpAtNet (22) 2024 RGB 3.8M - 93.33% - - -
(36) * 2023 RGB 5.34M - 98.15% - - 89.20%
LHGR-Net (218) 2023 RGB 1.81M - 93.15% - - 98.57%
2DPSTPP-Net (89) 2023 RGB 19.01M - - 100% 99.61% -
DRCAM (171) * 2023 Depth - 93.44% - - - -
EDenseNet (189) 2021 RGB 7.77M 99.99% - - 99.30% -
@@ 5.1.1 Spatiotemporal
In (200), the authors introduced a custom network, (2+1)D-SLR, composed of five stacked (2+1)D convolutional blocks (i.e., a 2D convolution followed by a 1D convolution). This structure lowers computational cost compared to traditional 3D convolutions, making the model more practical for deployment. To enhance temporal modeling, authors in (21) proposed a trainable key frame selection method based on error correction. VGG16 features are used to identify key frames, which are then preprocessed and fed into a custom 3DCNN for classification.
In (203), a dual-stage multi-stream pipeline for CSL recognition was developed. An adapted EfficientDet model—enhanced with cross-level Bi-FPN connections and a dual-channel spatial attention module (DCSAM)—detects both hands. The full image and cropped hand patches are passed through three parallel 3D-ResNet-18 models (following the P3D scheme), with their features fused and classified via an MLP. Authors in (70) introduced TS3C-Net, a modified 3D ConvNeXt that adds an extra 3×3×3 convolutional branch to each block, enhancing spatiotemporal representation. Finally, (191) addressed limitations in fixed-pipeline designs by combining features from both ResC3D and ConvLSTM using the Selective Spatiotemporal (SeST) module. The fused features are passed through fully connected and softmax layers for classification.
@@ 5.1.2 Spatial + Temporal
In (72), a method tailored for HCI tasks using only depth data was proposed. The method combined Inception-v3 that extracted spatial features with an LSTM for temporal modeling. For ISLR, authors in (196) introduced a pipeline to recognize Indian Sign Language (ISL) words used by deaf farmers. They combined GoogleNet for spatial encoding with a BiLSTM for sequence modeling.
To reduce inference time, a lightweight ISLR model, called TIM-SLR has been proposed in (201). The model uses a 2DCNN for per-frame feature extraction, followed by a Time Interaction Module (TIM)—a zero-parameter unit that fuses frame features into a compact representation. Finally, the Swin Transformer presented in (43) is a two-stage transformer-based model that extracts spatial features per frame, while a causally masked transformer processes the sequence, ensuring that only past and present frame features contribute at each step.
@@ 5.1.3 Short-Term Spatiotemporal + Long-Term Temporal
EgoFormer (155) is a model for egocentric dynamic gesture recognition optimized for edge devices, which has been introduced in the context of autonomous driving. EgoFormer combines a custom 3DCNN with a transformer module and an MLP for final classification. For long-range gesture-based UGV control, authors in (19) proposed a pipeline where a lightweight 2DCNN extracts per-frame features, which are used to select key frames. After cropping the human region, the selected frames are passed through a SlowFast network, followed by a transformer and an MLP.
To address data scarcity in sign language understanding, (23) proposed a zero-shot learning framework. Sign videos are represented using BERT-encoded textual descriptions combined with hand-crafted attributes. This representation guides a dual-stream backbone composed of short-term spatiotemporal modules (e.g., I3D) and long-term models (e.g., BiLSTM).
@@ 5.1.4 Other
Some appearance-based methods fall outside the previously discussed categories but offer valuable contributions, especially in Sign Language Recogition (SLR). For example, the method proposed in (126) for Indian SLR (ISLR) compresses an entire frame sequence into a single image, which is then processed by a MobileNet. A second MobileNet extracts per-frame features, which are passed through a two-layer LSTM. The outputs of both streams are concatenated and refined using a custom shallow 2DCNN. A different approach was presented in (104), where the authors introduced a dual-stream ISLR architecture combining Inception V4 with a novel convolutional module called stem.
In the context of pose-based ISLR, (123) proposed a dual-stream model with keyframe selection. A 2DCNN and attention mechanism filter out less informative frames. The remaining keyframes are processed by two branches: the Hand stream, which crops and analyzes hand regions with attention blocks, and the Full stream, which processes the entire frame. The fused features are used for final classification. To improve signer-independence and reduce overfitting, the SwC GR-MMixer (122), another lightweight model based on MLP-Mixer, incorporates GRU units for temporal modeling and a shifted window concatenation strategy to enhance sequence representation.